OpenStudy (anonymous):

I got 1 and 11/12 am I right? Evaluate. 5/2/3 - 2/1/4 x 1/2/3

OpenStudy (anonymous):

I got 3/4

OpenStudy (anonymous):

5 25/36

OpenStudy (anonymous):

1 and 11/ 12 and 5 and 25/36 are both an option ahhh! lol

OpenStudy (anonymous):

trust me lol

OpenStudy (anonymous):

I think the only way this can be done is to follow order of operations

OpenStudy (anonymous):

I'm confused on to what the equation is. Is this it? \[5\frac{ 2 }{ 3 } - 2\frac{ 1 }{ 4 } \times 1\frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

yeah altcode

OpenStudy (anonymous):

ohh crap ! @VeritasVosLiberabit

OpenStudy (anonymous):

let me double check now @Lindsay_552

OpenStudy (anonymous):

lol k so many comments at once !

OpenStudy (anonymous):

ah so there was a notation problem. Very important for this particular problem. if @AltCode is right this is much simpler

OpenStudy (anonymous):

\[\frac{ 10 }{ 3 }-(\frac{ 1 }{ 3 })=\frac{ 9 }{3 }=3\]

OpenStudy (anonymous):

Well sorry, I wasn't sure. That's jus what U got but im a Work in progress in math lol

OpenStudy (anonymous):

confussing problem lol

OpenStudy (anonymous):

If @AltCode has the problem written correctly I have the correct answer @Lindsay_552

OpenStudy (anonymous):

I met I got Lol I guess I cant type either lol

OpenStudy (anonymous):

yeah thats how it's wrote

OpenStudy (anonymous):

@VeritasVosLiberabit

OpenStudy (anonymous):

alright then check my reply above. The answer is 3